A jeepney was traveling with a velocity of 20 m/s when the driver stepped on the brake 

pedal. The jeepney traveled 4m away until it came to a full stop. How long did it take the 

jeepney to stop after the brake was applied?


1
Expert's answer
2021-06-13T07:08:55-0400
"d=\dfrac{1}{2}(v_0+v)t,""t=\dfrac{2d}{(v_0+v)},""t=\dfrac{2\cdot4\ m}{(20\ \dfrac{m}{s}+0)}=0.4\ s."